OpenStudy (anonymous):

Find the sum of the series 5+55+555+... to n terms, let f(n) be Sn, then verify that S2=60 when n=2 using f(n), and lastly, establish that 81/x f(n) is always a multiple of 81

OpenStudy (chaise):

Do you know the equation for the nth term? You're right, these are pretty hard.

OpenStudy (chaise):

a = 5 tn=a(n-1)+5*10^(n-1) Does this seem right?

OpenStudy (anonymous):

ahhhmmm, I've worked on it for so long, but there seems to be no progress

OpenStudy (anonymous):

I dont know....

OpenStudy (anonymous):

first term is 5*1 2nd= 5*12 3rd=5*123 4th=5*1234 Do you see the rule?

OpenStudy (anonymous):

still I do not see how to change this into direct equation

OpenStudy (anonymous):

yeah, maths HL in the IB diploma is unbelievably hard

OpenStudy (anonymous):

what is HL and IB?

OpenStudy (anonymous):

I am thinking of writing it up as a sum

OpenStudy (anonymous):

\[5*\sum_{1}^{n} n10^{n-1}\]

OpenStudy (anonymous):

Do you think it is right? :-)

OpenStudy (anonymous):

I will post it again so someone can check it, ok?

OpenStudy (anonymous):

HL is for higher level and IB is for International Bacchalerette

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

the answer is 5/81(10^(n+1)−9n−10)

OpenStudy (anonymous):

check my topic if you need the calculation

OpenStudy (anonymous):

thank you so much! but what do you mean by check my topics?

OpenStudy (anonymous):

I reposted ur question

OpenStudy (anonymous):

loved this problem btw! So thanks for posting

OpenStudy (anonymous):

no, thank YOU for helping me with this! my maths teacher gaved these questions to me and told me that it took her a long time to do them too :-/